# Know Your Neighborhood: General and Zero-Shot Capable Binary Function Search Powered by Call Graphlets

### v1

@@ Abstract
Binary code similarity detection is an important problem with applications in areas like malware analysis, vulnerability research and plagiarism detection. This paper proposes a novel graph neural network architecture combined with a novel graph data representation called call graphlets. A call graphlet encodes the neighborhood around each function in a binary executable, capturing the local and global context through a series of statistical features. A specialized graph neural network model is then designed to operate on this graph representation, learning to map it to a feature vector that encodes semantic code similarities using deep metric learning.
The proposed approach is evaluated across four distinct datasets covering different architectures, compiler toolchains, and optimization levels. Experimental results demonstrate that the combination of call graphlets and the novel graph neural network architecture achieves state-of-the-art performance compared to baseline techniques across cross- architecture, mono-architecture and zero shot tasks. In addition, our proposed approach also performs well when evaluated against an out-of-domain function inlining task. Overall, the work provides a general and effective graph neural network-based solution for conducting binary code similarity detection.
@@ 1 Introduction
Binary code similarity detection (BCSD) aims to find compiled functions that are similar to a given query function. It has a wide range of applications, from identifying vulnerabilities in firmware to analyzing code reuse across malware variants. BCSD can enhance security assessments of closed-source software by accurately identifying the components used, such as libraries embedded within statically linked executables and firmware binaries.
@@
Existing approaches predominantly focus on function-level artifacts like control flow graphs, disassembly, or intermediate representations, neglecting the potential of call graph features. Additionally, some approaches rely on the assumption of dynamically linked standard library functions such as those from libc are easily recoverable, which does not hold for embedded devices and malware where static linking is common, restricting their applicability across different software ecosystems.
Furthermore, the evaluation methodologies employed in binary function search often lack rigor. Existing approaches predominantly rely on "in-domain" evaluation data, utilizing subsets of functions from the training binaries themselves. This practice introduces the risk of data leakage, potentially inflating reported performance metrics. More crucially, it fails to assess the true generalization capability of the approach when faced with previously unseen functions, a scenario referred to as "out-of-domain" data. Comprehensive evaluation necessitates testing on entirely separate datasets to obtain an accurate understanding of real-world performance and generalizability. This is an area we tackle head on with evaluation across four distinct datasets.
This paper proposes a novel approach that aims to exploit the additional information available when including call graph features and addresses the limitations of existing methods. It employs a comprehensive evaluation methodology, including both in-domain and out-of-domain assessments across four large datasets, to ensure robust and generalizable findings. The proposed approach does not rely on the identification of standard library functionality, making it applicable across a broader range of binary types and scenarios.
@@ 1.1 Key Contributions
    We present a simple, yet powerful Graph Neural Network (GNN) trained using deep metric learning which is capable performing well in and out of domain across a number of tasks.
@@
    We present a robust evaluation using previous research datasets of our approach as well as compare against a range of benchmarks.
@@
    We present conduct an ablation study to investigate how our graph neural network (GNN) design choices contribute to our performance.
@@ 1.2 Paper Overview
The remainder of this paper is organized as follows: We begin by reviewing the relevant prior work to establish a foundation for our contributions. Next, we present a detailed description of the methodology used in this study. This includes explaining how a call graphlet is constructed, the features used for nodes and edges, the Graph Neural Network (GNN) architectural choices, and the datasets employed. Following this, we present a comprehensive section on the experiments conducted to evaluate our approach. We start with an evaluation in a cross-architecture context, followed by experiments on the model’s performance in a mono-architecture setting (x86-64 only), cross-architecture with function inlining setting, zero-shot vulnerability search setting, and then conclude with an ablation study. Finally, we conclude the paper by exploring some fundamental limitations of our approaches and providing our concluding remarks.
@@ 2 Related Work
The problem of identifying similar functions across different binary architectures has been tackled using various methods in the existing literature. These methods can generally be classified into three main categories: sequence-based approaches and graph-based approaches, or a combination of both. In this section, we start by briefly reviewing the sequence-based approaches. Then, we provide a more detailed review of the graph-based literature. Finally, we will discuss alternative methods and highlight the limitations of previous works in this area.
2.1 Sequence Based Approaches
Sequence based approaches use either assembly, pseudo-code or intermediate representations extracted from binaries for each function to generate representations suitable for comparison. These sorts of approaches have been applied to individual assembly instructions[13], sequences of assembly instructions[23], sequences of intermediate representations[4] and even execution sequences [18].
All approaches share the common objective of seeking to learn useful representations of input by adapting approaches developed in the natural language processing (NLP) literature. This is typically done via a two stage process of pre-training and then fine-tuning. Researchers have used a range of different pre-training tasks ranging from standard approaches such as Masked Language Modelling (MLM) to domain specific tasks such as jump target prediction [23] and instruction data flow relationship prediction [13]. Once pre-trained, the models are then fine-tuned for a given down stream task such as binary code similarity detection (BCSD) or N-day vulnerability detection typically using contrastive methods such as deep metric learning.
@@ 2.2 Graph Based Approaches
All of these approaches however miss an opportunity to use broader information related to a single function such as type information and calling relationships and instead focus on data contained within the a single function boundary. Graph based approaches on the other hand, look to exploit the inherent data structure within compiled binaries such as call graphs and control flow graphs (CFG). These data structures are then paired with graph neural network approaches to create learnt representations that perform well at downstream tasks such as BCSD.
XBA [11] is a graph neural networks approach which works on a basic block level for a given function as well as incorporating key external information such as string literals and external function calls. This novel representation is then fed into a GCN based GNN to generate cross-architecture and cross-platform embeddings. PDM [17] is another example of a graph neural network which instead works on a combined control flow and data flow representation named ACFG+. This representation is then used alongside a Capsule GNN to generate embeddings.
CFG2VEC [25] is an example of a more complicated approach to ours and was used as inspiration for our approach. CFG2VEC is a hierarchical method whereby a single input is a binaries entire call graph and each node is a P-code Based control flow graph. The P-Code based CFG representation is then fed into a graph neural network to create node-embeddings for the higher level call graph. These node embeddings are then subsequently used to support binary code similarity detection tasks. PSS [2] propose an interesting graph spectra approach which operates primarily on the call graphs structure. This approach is used for whole binary identification so can be considered tangential to the other papers within this section but has been included primarily because it was also a large the main inspiration for this work. HermesSim [8] also used a P-Code representation of a target function but instead creates a novel graph representation where nodes are P-code operations or memory locations and the edges are relationships such as read and writes. This is again fed into a GNN approach to create function level representations that can be used for BCSD tasks. HermesSim can be considered the current state of the art in this area having reported impressive results recently.
@@ 2.3 Alternative Methods
In addition to sequence, graph and combinatorial methods, some researchers have sort to adopt alternative methods. BinFinder [19] is an example of this were instead of using sequences, a series of statistical count features were used based on a VEX IR lifted representation and counts of glibc calls. This approach has one major drawback and that is the reliance on glibc functions being recoverable and external. This has potential to limit its usage in cases such as statically linked binaries or blob firmware. The approach however does perform very well and is one our a key baselines.
@@ 2.4 Limitations of Previous Work
There are several limitations of the works mentioned above. Firstly, several of the approaches [19, 11] rely on recoverable function names to build fundamental parts of there feature space. This is error-prone and something that may not always be possible in all settings such as when you are faced with obfuscated or statically linked binaries. Secondly, most existing approaches [17, 13, 4, 23, 18] analyze individual functions in isolation without considering the broader relationships and interactions between functions within the software binary. This potentially overlooks valuable information that is inherent in the overall structure and interconnections present in the binary.
@@ 3.1 Call Graphlet Neighborhoods
Our work utilizes a novel data format we call call graphlets (visually depicted Figure 1). This approach was inspired by the work described in PSS [2] but instead of operating with binary level call graphs, we opt to work with function level ones. A call graphlet is a weighted directed graph made up of a tuple $G=V,E,W$ whereby: $V$ represents a set of functions, $E$ represents a set of calling relationship in the format of $(x,y)\;|\;x\neq y$ and $W$ represent the weights for a given edge. A target function’s call graphlet neighborhood includes a node for the target function itself, any functions that call the target function (callers), any functions the target function calls (callees) and any functions that are called by the target functions callees (callees of callees). We excluded callers of callers from our call graphlets to prevent creating excessively large call graphlets, particularly for widely used functions like utilities and standard libraries, and because we felt that callee relationships better characterize a function. The weights for all edges correspond to there respective edge betweenness value that is calculated at a call graphlet level rather than a global call graph level.
The reasons for working at a call graphlet level are twofold. Firstly, recent work suggests that features derived from the call graph are the most robust to changes in compilation toolchains [10]. We however viewed using global call graphs as too coarse and instead formulated a more fine grained data representation with the aim of developing a high performing function-level BCSD approach. Secondly, we wanted to create a data representation that took the calling behaviour between a function, its callers and callees into account. We view this relationship as something that will likely be similar within software compiled from the same source code regardless of compilation variations. This provides a means to focus on a higher level representation rather than solely focusing on function level differences or changes such as patches or refactoring.
@@ 3.2.1 Node Features
Each of the functions within a given call graphlet neighborhood have the six straightforward features derived from function level metadata recovered during the disassembly process. Features such as in-degree (number of callers), out-degree (number of callees) and total number of edges provide global information whilst features such as total number of instructions, number of arguments and number of local variables provide local information. The hypothesis is that this combination of local and global level information provides the model the ability to discern between a range of different functions. For example, a complex function that is infrequently called may have a low in-degree but a high number of instructions, out-degree and number of local variables whereas a wrapper function for a commonly used library function such as memcpy may have a low instruction count and out-degree but a very high in-degree suggesting it is called frequently. The full feature list can be found in Table 1.
@@
Table 1: Function/Node Level Features
@@ 3.2.2 Edge Weights
The edge betweenness is calculated at a call graphlet level instead of a global call graph level because we want to quantify the control each edge exerts over information flow within the call graphlet. Edges with high betweenness centrality can be viewed as critical pathways or bottlenecks. In our particular case, we viewed edge betweenness centrality as a useful addition because it provides a means of identifying functions that exhibit high call fan-out (i.e., a large number of callees) and participate in critical calling relationships with ancestor functions in the call hierarchy. This additional information is hypothesized to improve the model’s performance by capturing the interplay between a function’s calling behavior and its position within the program’s overall call structure.
@@ 3.3.1 Training Dataset
It is worth noting that the processed call-graphlet dataset created from Dataset-1 using bin2ml is significantly larger than what is reported in [15] and [8]. This is because our data representation allows us to relax the filtering requirements such as number of instruction or basic blocks within a function. Since our method uses a functions neighborhood, we can include smaller functions as they will be augmented by their neighbors who may not fit the filtering requirements.
Our data representation also likely results in more varied examples of each function because, due to architectural, compiler or optimization differences, the variations present within a functions neighborhood are included as well. This is especially interesting because even when the node features of a function are duplicates, each call graphlet may be different and therefore, continue to be included within our dataset. With this being said, in order to facilitate robust and fair comparison, we firstly subset the data into the same binary-level splits described in [15] and then use random sampling to create identically sized samples from within the binaries within the split. This methodology is replicated to generate the corresponding Dataset-1 test set split as well.
@@ 3.3.2 Evaluation Datasets
In order to robustly evaluate our approach, we use four open source benchmark datasets and focus exclusively on out-of-domain evaluation. The reason for focusing exclusively on out-of-domain performance is that it provides experimental evidence of our approaches performs well when given a collection of binaries it has not already seen as well as being the most comparable to a real-world setting.
Firstly, we use the test set from Dataset-1. This dataset includes a subset binaries which encompasses the Z3 and Nmap projects, all of which are not present within the train dataset. This dataset contains 522003 functions which is identical to the size used within [8] and [15].
@@
Thirdly, we use BinKit NoInline dataset presented within [10]. This dataset contains 51 unique packages/libraries across 8 architectures (x86, x64, ARM32, ARM64, MIPS32, MIPS64, MIPS32EB and MIPS64EB), 9 different versions of compiler (five unique versions of GCC, four unique versions of Clang) as well as five different optimization levels (O0, O1, O2, O3, Os). All of the binaries have also been compiled with function inlining turned off. This results in a total of 67,680 binaries. This dataset does share some common software projects to Dataset-2 but does not share any with the training dataset. This dataset introduces two previously unseen architecture variants, additional compilers as well as an additional optimization level.
And finally, we use the BinKit Normal dataset also presented within [10]. This is an identical dataset to the Binkit NoInline dataset described above but with one fundamental difference - The compilers are able to inline functions. Function inlining has been referenced as a significant weakness across a number of binary function similarity approaches [24, 6] and therefore this dataset allows us to tackle this area head on and generate empirical results of how well our approach fairs when faced with function inlining.

### v2

@@ Abstract
Binary code similarity detection is an important problem with applications in areas such as malware analysis, vulnerability research and license violation detection. This paper proposes a novel graph neural network architecture combined with a novel graph data representation called call graphlets. A call graphlet encodes the neighborhood around each function in a binary executable, capturing the local and global context through a series of statistical features. A specialized graph neural network model operates on this graph representation, learning to map it to a feature vector that encodes semantic binary code similarities using deep-metric learning.
The proposed approach is evaluated across five distinct datasets covering different architectures, compiler tool chains, and optimization levels. Experimental results show that the combination of call graphlets and the novel graph neural network architecture achieves comparable or state-of-the-art performance compared to baseline techniques across cross- architecture, mono-architecture and zero shot tasks. In addition, our proposed approach also performs well when evaluated against an out-of-domain function inlining task. The work provides a general and effective graph neural network-based solution for conducting binary code similarity detection.
@@ 1 Introduction
Binary code similarity detection (BCSD) aims to find compiled functions that are similar to a given compiled query function. It has a wide range of applications, from identifying vulnerabilities in firmware to analyzing code reuse across malware variants. BCSD can enhance security assessments of closed-source software by identifying the components used, such as libraries embedded within statically linked executables and firmware binaries.
@@
Existing approaches predominantly focus on function-level artifacts like control flow graphs, disassembly, or intermediate representations, neglecting the potential of call graph features. Additionally, some approaches rely on the assumption of dynamically linked standard library functions such as those from libc are easily recoverable. This does not hold for embedded devices and malware where static linking is common, restricting their applicability across different software ecosystems.
Furthermore, the evaluation methodologies employed in binary function search often lack rigor. Existing approaches predominantly rely on “in-domain” evaluation data, utilizing subsets of functions from the training binaries themselves. This practice introduces the risk of data leakage, whereby the model is evaluated on the same data in which it was trained leading to the potential of reporting inflated performance metrics. More crucially, it fails to assess the true generalization capability of the approach when faced with previously unseen functions, a scenario referred to as “out-of-domain” data. Comprehensive evaluation necessitates testing on entirely separate datasets to obtain an accurate understanding of real-world performance and generalizability. This is an area we tackle head on with evaluation across five distinct datasets.
This paper proposes a novel approach that aims to exploit the additional information available when including call graph features and addresses the limitations of existing methods. It employs a comprehensive evaluation methodology focusing on out-of-domain assessments across four large datasets, to ensure robust and generalizable findings. The proposed approach does not rely on the identification of standard library functionality, making it applicable across a broader range of binary types and scenarios.
@@ 1.1 Key Contributions
    We present a simple, yet powerful, Graph Neural Network (GNN) trained using deep-metric learning capable of performing well in and out of domain across a number of tasks.
@@
    We present a robust evaluation of our approach using a range of research dataset benchmarks and compare against both sequence and graph based state-of-the-art approaches.
@@
    We conduct an ablation study to investigate how our graph neural network (GNN) design choices contribute to our performance.
The data, models and associated experimental code is available: https://github.com/br0kej/kyn
@@ 1.2 Paper Overview
The remainder of this paper is organized as follows: We begin by reviewing the relevant prior work to establish a foundation for our contributions. Next, we present a detailed description of the methodology used in this study. This includes explaining how a call graphlet is constructed, the features used for nodes and edges, the Graph Neural Network (GNN) architectural choices, and the datasets employed. Following this, we present a comprehensive section on the experiments conducted to evaluate our approach. We start with an evaluation in a cross-architecture context, followed by experiments evaluating the model’s performance in a mono-architecture setting (x86-64 only), a cross-architecture with function inlining setting, a zero-shot vulnerability search setting, and then conclude with an ablation study. Finally, we conclude the paper by exploring some fundamental limitations of our approaches and providing our concluding remarks.
@@ 2 Related Work
The problem of identifying similar functions across different instruction set architectures has been tackled using various methods in the existing literature. These methods can generally be classified into two main categories: sequence-based approaches or graph-based approaches. In this section, we start by briefly reviewing the sequence-based approaches. Then, we provide a more detailed review of the graph-based literature. Finally, we briefly introduce a relevant alternative method and before then highlighting the limitations of previous works in this area.
2.1 Sequence-Based Approaches
Sequence-based approaches use either assembly, pseudocode or intermediate representations extracted from binaries for each function to generate representations suitable for similarity comparison. These sorts of approaches have been applied to individual assembly instructions[3], sequences of assembly instructions[4], sequences of intermediate representations[5] and even execution sequences [6].
All approaches share the common objective of seeking to learn useful representations of input by adapting approaches developed in the natural language processing (NLP) literature. This is typically done via a two stage process of pre-training and then fine-tuning. Researchers have used a range of different pre-training tasks ranging from standard approaches such as Masked Language Modelling (MLM) to domain specific tasks such as jump target prediction [4] and instruction data flow relationship prediction [3]. Once pre-trained, the models are then fine-tuned for a given down stream task such as binary code similarity detection (BCSD) or N-day vulnerability detection typically using contrastive methods such as deep-metric learning.
@@ 2.2 Graph Based Approaches
All of these approaches however miss an opportunity to use broader information related to a single function such as type information and calling relationships and instead focus on data contained within a single function boundary. Graph based approaches on the other hand, look to exploit the inherent structure within compiled binaries such as call graphs and control flow graphs (CFG). These data structures are then paired with graph neural network approaches to create learned representations that perform well at downstream tasks such as BCSD or N-day vulnerability detection.
XBA [7] is a graph neural networks approach which works on a basic block level for a given function as well as incorporating key external information such as string literals and external function calls. This novel representation is then fed into a GCN-based GNN to generate cross-architecture and cross-platform embeddings. PDM [8] is another example of a graph neural network which instead works on a combined control flow and data flow representation named ACFG+. This representation is then used alongside a Capsule GNN to generate embeddings.
CFG2VEC [9] is an example of a more complicated approach to ours and was used as inspiration for our approach. CFG2VEC is a hierarchical method whereby a single input is the call graph of the binary of interest and each node is a P-code Based control flow graph. The P-Code based CFG representation is then fed into a graph neural network to create node-embeddings for the higher level call graph. These node embeddings are then subsequently used to support binary code similarity detection tasks. PSS [2] propose an interesting graph spectra approach which operates primarily on the call graphs structure. This approach is used for whole binary identification so can be considered tangential to the other papers within this section but has been included primarily because it was also a large the main inspiration for this work. HermesSim [10] also used a P-Code representation of a target function but instead creates a novel graph representation where nodes are P-code operations or memory locations and the edges are relationships such as read and writes. This is again fed into a GNN approach to create function level representations that can be used for BCSD tasks. HermesSim can be considered the current state of the art in this area having reported impressive results recently and is the primary baseline we use within this work.
@@ 2.3 Alternative Methods
In addition to sequence, graph and combinatorial methods, some researchers have sought to adopt alternative methods. BinFinder [11] is an example of this were instead of using sequences, a series of statistical count features that are calculated from the VEX IR lifted representation of a function alongside counts of function calls such as those to glibc. This approach has one major drawback and that is the reliance on glibc functions being recoverable and external. This limits its usage in cases such as statically linked binaries or blob firmware. The approach however does perform very well and is one our key baselines.
@@ 2.4 Limitations of Previous Work
There are several limitations of the works mentioned above. Firstly, several of the approaches [11, 7] rely on recoverable function names to build fundamental parts of their feature space. This is error-prone and something that may not always be possible in all settings such as when you are faced with obfuscated or statically linked binaries. Secondly, most existing approaches [8, 3, 5, 4, 6] analyze individual functions in isolation without considering the broader relationships and interactions between functions within the software binary. This potentially overlooks valuable information that is inherent in the overall structure and interconnections present in the binary.
@@ 3.1 Call Graphlet Neighborhoods
To overcome these limitations, our work utilizes a novel data format we call call graphlets which is shown in Figure 1. This approach was inspired by the work described in PSS [2] but instead of operating with binary-level call graphs, we opt to work with function-level ones which provide a balance of local and global information.
A call graphlet is a weighted directed graph made up of a tuple $G=(V,E,W)$ where $V$ represents a set of functions, $E$ represents a set of calling relationship in the format of $(x,y)\;|\;x\neq y$ and $W$ represent the weights for a given edge. A target function’s call graphlet neighborhood includes a node for the target function itself, any functions that call the target function (callers), any functions the target function calls (callees) and any functions that are called by the target functions callees (callees of callees). We excluded callers of callers from our call graphlets to prevent creating excessively large call graphlets, particularly for widely used functions like utilities and standard libraries, and because we felt that callee relationships better characterize a function. The weights for all edges correspond to there respective edge betweenness which is calculated at a call graphlet level rather than a global call graph level.
The reasons for working at a call-graphlet level are twofold. Firstly, recent work suggests that features derived from the call graph are the most robust to changes in compilation tool chains [1]. We however viewed using global call graphs as too coarse and instead formulated a more fine-grained data representation with the aim of developing a high performing function-level BCSD approach. Secondly, we wanted to create a data representation that took the calling behaviour between a function, its callers and callees into account. We view this relationship as something that will likely be similar within software compiled from the same source code regardless of compilation variations. This provides a means to focus on a higher level representation rather than solely focusing on function level differences which can be effected by not only toolchain choice but also changes such as patches or refactoring.
@@ 3.2.1 Node Features
Each of the functions within a given call graphlet neighborhood have six straightforward features derived from function level metadata recovered during the disassembly process. Features such as in-degree (number of callers), out-degree (number of callees) and total number of edges provide global information whilst features such as total number of instructions, number of arguments and number of local variables provide local information. The hypothesis is that this combination of local and global level information provides the model the ability to discern between a range of different functions. For example, a complex function that is infrequently called may have a low in-degree but a high number of instructions, out-degree and number of local variables whereas a wrapper function for a commonly used library function such as memcpy may have a low instruction count and out-degree but a very high in-degree. The full feature list can be found in Table I.
@@
TABLE I: Function/Node Level Features
@@ 3.2.2 Edge Weights
The edge betweenness is calculated at a call graphlet level instead of a global call graph level because we want to quantify the control each edge exerts over information flow within the call graphlet. Edges with high betweenness centrality can be viewed as critical pathways or bottlenecks. In our particular case, we viewed edge betweenness centrality as a useful addition because it provides a means of identifying functions that exhibit high call fan-out (i.e., a large number of callees) and participate in critical calling relationships with ancestor functions in the call hierarchy. This additional information is hypothesized to improve the model’s performance by capturing the interplay between a function’s calling behavior and its position within call graphlet’s overall structure.
@@ 3.3.1 Training Dataset
It is worth noting that the processed call-graphlet dataset created from Dataset-1 using bin2ml[13] is significantly larger than what is reported in [12] and [10]. This is because our data representation allows us to relax the filtering requirements such as number of instruction or basic blocks within a function. Since our method uses a function’s neighborhood, we can include smaller functions as they will be augmented by their neighbors who may not fit the filtering requirements.
Our data representation also likely results in more varied examples of each function because, due to architectural, compiler or optimization differences, the variations present within a function’s neighborhood are included as well. This is especially interesting because even when the node features of a function are duplicates, each call graphlet may be different and therefore, continue to be included within our dataset. With this being said, in order to facilitate robust and fair comparison, we firstly subset the data into the same binary-level splits described in [12] and then use random sampling to create an identically sized sample from within the binaries previously subset. This methodology is used to generate the both the training and tests sets for Dataset-1.
@@ 3.3.2 Evaluation Datasets
In order to robustly evaluate our approach, we use five open source benchmark datasets and focus exclusively on out-of-domain evaluation. The reason for focusing exclusively on out-of-domain performance is that it provides experimental evidence that our approach performs well when faced with a collection of binaries it has not seen before as well as being the most comparable to a real-world setting.
Firstly, we use the test set from Dataset-1. This dataset is a collection of binaries which encompasses the Z3 and Nmap projects, all of which are not present within the train dataset. This dataset contains 522003 functions which is identical to the size used within [10] and [12].
@@
Thirdly, we use BinKit NoInline dataset presented within [1]. This dataset contains 51 unique packages/libraries across 8 architectures (x86, x64, ARM32, ARM64, MIPS32, MIPS64, MIPS32EB and MIPS64EB), 9 different versions of compiler (five unique versions of GCC, four unique versions of Clang) as well as five different optimization levels (O0, O1, O2, O3, Os). All the binaries have also been compiled with function inlining turned off. This results in a total of 67,680 binaries. This dataset does share some common software projects to Dataset-2 but does not share any with the training dataset. The dataset introduces two previously unseen architecture variants, additional compilers as well as an additional optimization level.
Fourthly, we use the BinKit Normal dataset also presented within [1]. This is an identical dataset to the Binkit NoInline dataset described above but with one fundamental difference - The compilers are able to inline functions. Function inlining has been referenced as a significant weakness across a number of binary function similarity approaches [14, 15] and therefore this dataset allows us to tackle this area head on and generate empirical results of how well our approach fairs when faced with function inlining.
Finally, we use the Dataset-Vulnerability dataset presented as part of [12]. This dataset is a collection of OpenSSL libraries compiled for a range of architectures alongside two examples of the identical library but instead taken from a real device, one which is arm32 and another which is mips32. This forms the basis for a small scale N-day vulnerability benchmark.